In a clinical​ trial, 23 out of 870 patients taking a prescription drug daily complained of flulike symptoms. Suppose that it is known that 2.4​% of patients taking competing drugs complain of flulike symptoms. Is there sufficient evidence to conclude that more than 2.4​% of this​ drug's users experience flulike symptoms as a side effect at the α=0.05 level of​ significance?

Because np0(1−p0)=__?__

<

>

=

​10, the sample size is

less than

greater than

​5% of the population​ size, and the sample

can be reasonably assumed to be random,

is given to be random,

cannot be reasonably assumed to be random,

is given to not be random,

the requirements for testing the hypothesis

are not

are

satisfied.

​(Round to one decimal place as​ needed.)
Answer #1

np0(1−p0) = 20.3788 > 10 the sample size is greater than ​5% of the population​ size, and the sample can be reasonably assumed to be random, is given to be random the requirements for testing the hypothesis are not satisfied.

Z = 0.4 ( one decimal place )

P value = 0.4 ( one decimal place )

Calculations :

Z critical = 1.645 > Z stat

Do not reject null hypothesis

there is no sufficient evidence to conclude that more than 2.4​% of this​ drug's users experience flulike symptoms as a side effect at the α=0.05 level of​ significance